Assume that calendar year 2017 ended a few weeks ago.

Prepare Soaring Company’s Balance Sheet at December 31, 2017 and its Income and Expense Statement for calendar year 2017 from the information in the Table below.

The Balance Sheet and Income and Expense Statements will provide the basis for an elementary financial analysis of the Soaring Company.

The income tax rate on capital gains, corporate income and recaptured depreciation is 40%.

Reference Description Amount

1. Accounts payable by January 23, 2018                                                                                             $233,000

2. Accounts receivable by February 28, 2018.                                                                                      $255,000

3. Unpaid wages payable before January 31, 2018.                                                                             $25,000

4. Cash $20,000

5. Common Share Capital on December 31, 2017. $577,000

6. Goods purchased for resale in 2017. $5,665,000

7. Cumulative depreciation of plant on December 31, 2017. $760,000

8. Cumulative depreciation of equipment on December 31, 2017. $649,000

9. Plant depreciation in 2017. $45,000

10. Equipment depreciation in 2017. $33,000

11. Historical cost of equipment purchases $789,000

12. 2017 federal income taxes payable by April 20, 2018. $15,000

13. Interest paid on loans paid in 2017. $17,600

14. Inventory of finished goods on December 31, 2017 $113,200

15. Inventory of unfinished goods on December 31, 2017 $55,900

16. Land $450,000

17. 10-year bonds issued by Soaring Company on June 15, 2012. $100,000

18. Short-term provincial bonds purchased by Soaring Company in December2017 $25,000

19. Mortgage due in December 2020 $380,000

20. Miscellaneous operating expenses in 2017 $47,800

21. Cumulative cost of plant (building purchases) over the years $1,450,000

22. Prepaid expenses (computer services ….) in 2017 $900

23. Soaring Company total sales in 2017 $6,258,000

24. Retained earnings (after dividends) on December 31, 2017. $555,000

25. Warehouse rental expense in 2017 $34,000

26. Wages and salaries paid in 2017. $166,000

27. Water, hydro, heating and property taxes in 2017. $33,900

28. Interest income in 2017 from financial investments. $6,970

29. Market price of Soaring Company common shares on December 31, 2017. $20

30. Dividends per common share declared and paid on December 21, 2017. $0.65

31. Number of Soaring Company common shares on December 31, 2017. $60,000

Question

1.a) Soaring Company’s interest expense coverage (1 st decimal; no rounding) is a) 8.8; b) 9.1; c) 9.5; d) 11.6.

b) Soaring Company’s earnings per share (2 nd decimal; no rounding) is a) $2.22; b) $2.41; c) $2.53; d) $3.61.

c) Soaring Company’s price-earnings ratio (1 st decimal; no rounding) is a) 8.6; b) 8.9; c) 9.4; d) 10.9.
Answer #1

1 Balance Sheet 2 Assets 3 Current Assets 4 Accounts Receivable 5 Cash 6 Invento Income and Expense Statement Particulars Sal

1.a) Interest Coverage Ratio = EBIT / Interest Expense = 233,300/17600 = 13.2 [13.2557]

1.b) EPS = Earnings After Tax / Number of Shares = 133,602 / 60,000 = 2.22 [2.2267]

1.c) Price Earning Ratio = Market Price / EPS = 20 / 2.22 = 8.9 [8.9819]

Good luck